OpenStudy (anonymous):

If cot thet= 3/4 and the terminal point determined by theta is in quadrant 3, then: a) tan theta =4/3 b) cos theta = -3/5 c) sin theta =3/5 d) csc theta= -5/3

OpenStudy (anonymous):

To find cos x, sin x, and csc x, use these trig identities: 1 + cot^2 x = 1/sin^2 x = csc^2 x 1 + tan^2 x = 1/cos^2 x